Systems Staff Meeting Notes 5/1/00

Systems Staff Meeting Notes
May 1, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. Bobbie's dog, Katie, is feeling better - she played with her toy star yesterday!
  2. Logo#2 was picked for the Systems T-shirts.
Ongoing Systems Updates
Orders
  1. A three-foot wide printer is on order for Documents. It will be able to print pages up to 300 feet long. The Libraries did not want to buy Census maps so this printer will be used to print them on demand - at $2.00-$6.00 a piece.
  2. Two or three laptops have been ordered.
  3. PDA's - Personal Digital Assistants - a few have been ordered: Palm Pilots, Windows Pockets PC's. Initially, they will be destined for the top Library Personnel - eventually, they will probably trickle down to other departments.
Equipment Received/Installed/Repaired
  1. Gerald Langford reported that the very old (1991) IBM LaserPrinter E that they use for printing labels in Resource Services has died. We will give them an Okidata that we have on the shelf, but they will probably want to order a new printer.
Liaisons
  1. Liaison meeting scheduled for this Thursday, May 4.
Networking
  1. IPX problems continue - the fast machines in 206 MSL are still refusing to speak to IPX. Bill & Will checked on the problem on Friday without solution. Will try a fast laptop to eliminate the chance that it's a configuration problem.
  2. Novell CD conversion to NT - NT10 is alive& rebuilt, and a new version of the CD software has been downloaded. Will wait for an appropriate time to test.
  3. Wireless - Brad (Facilities) is still hung on the idea that the conduit does not need to run back to the closet from the stacks. He has suggested a cable tie to the existing conduit. Will will get Judy to take a look at it to determine what is necessary. Need to check on the legality - it may require permits.
  4. West wiring - projected completion date of mid-May still holds.
  5. Journalism - started the project this morning, need to do a bit more testing. There was no light when the switch was connected to the new fiber patch. Bill reminds to check whether the feed is 100M or if it is still 10M.
  6. West - Business Services - no news on the room for servers.
Software Projects Status
  1. Film Log - will demo again this Thursday, hopefully this will rap it up.
  2. Library Statistics - the Health Center ban on IE use seems to come from Bill Koontz, not the Health Center. Catalyst functions require it.
  3. Ying is researching the add-ins on Outlook/Exchange for a PAULA type function - so far, no luck.
DP Coordinator
  1. Laptop checkout on Exchange - Debra has been testing the setup by Suzy. Seems like it will work out just fine. Debra is working on a listing of configurations for each laptop. In reviewing each laptop, some problems were found: one needs a Xircom modem/networking card to be ordered for it, one needs the second disk of Win'98 loaded & one needs to be upgraded to Win'98.
  2. Students out until Summer Semester, which starts on May 15th.
Calendar
  1. Bobbie out May 4 & 5.
  2. Hilton out May 22.
  3. Debra out May 26.
  4. Memorial Day - Monday, May 29.
  5. Winston will be out June 19-23.
  6. Bobbie will be out July 10-21.
  7. Will is scheduled to be out July 24-28.


Back to the Systems Homepage

Last updated May 8, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu